Find the monthly payments on a $2400 loan made for 18 months at 12% interest rate?

Please assist in solution and how to find solution

Answer by stanbon(48554)

Find the monthly payments on a $2400 loan made for 18 months at 12% interest rate?
--------
Loan Balance = (original amount)(1+i)^n - [P/i][(1+i)^n - 1]
where i = periodic interest rate; n=# of periods; P = periodic payment
------------
You want the loan balance to be zero
---------------
Your problem:
i=0.12/18= 0.0067
0 = 2400(1.0067)^18 - [P/0.0067][1.0067^18 - 1]
0 = 2706.53 -[P/0.0067][0.1277199]

P/0.0067 = 21191.13779
Monthly Payment = $141.98
========
Cheers,
Stan H.